OpenStudy (anonymous):

Complete the square Vsquared + 10v -21 = 0

OpenStudy (anonymous):

take half the middle coefficent........

OpenStudy (anonymous):

so 5v

OpenStudy (anonymous):

just 5

OpenStudy (anonymous):

square it

OpenStudy (anonymous):

then add and subtract that value from the left side of the equation, should look something like this: v^2+10v+25-25-21=0 (v+5)^2-46=0

OpenStudy (anonymous):

thank you